OpenStudy (anonymous):

Kenny mowed a lawn today to earn some extra money. He spent $1.40 on fuel and $3.25 on lunch. He charged $15 to mow the lawn. What was his profit for mowing the lawn?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (texaschic101):

what he charged - what he bought = profit 15 - (1.40 + 3.25) = profit can you figure it out now ?

OpenStudy (anonymous):

16.85

OpenStudy (texaschic101):

15 - (1.40 + 3.25) = profit 15 - 4.65 = profit 10.35 = profit

OpenStudy (anonymous):

yep thats what I got after I redid it thanks!

OpenStudy (texaschic101):

sure thing :)